# Rheo-SINDy: Finding a constitutive model from rheological data for complex fluids using sparse identification for nonlinear dynamics

### v2

@@ IV.2 FENE-P Dumbbell Model
Figure 7 shows (a) the total number of terms and (b) the error rate of the constitutive equation obtained by Rheo-SINDy with the STRidge and a-Lasso for the FENE-P dumbbell model. Similar to the previous case, the a-Lasso yields sparser solutions than the STRidge. Based on the number of terms and the error rates shown in Fig. 7, we chose several $\alpha$ values with a small number of terms and a low error rate. Figure 8 presents the equations obtained using the chosen $\alpha$. From Fig. 8, the equations predicted by the STRidge with $\alpha=1$ and the a-Lasso with $\alpha=1\times 10^{-4}$ are almost the same; conversely, the solutions for small $\alpha$ values significantly differ between the two methods. For the STRidge with $\alpha=1\times 10^{-3}$, the identified equations are close to the correct equations (cf. Eqs. (29)–(31)). Furthermore, the coefficient values for the terms obtained correctly are close to the correct values. For the a-Lasso with $\alpha=1\times 10^{-8}$, several coefficients for the correctly obtained terms, such as $\tau_{xx}$, $\tau_{xy}\kappa_{xy}$, and $\tau_{xx}\tau_{xy}\kappa_{xy}$ in the equation for $\dot{\tau}_{xx}$, are close to the exact values, but for other several terms, such as ${\rm tr}\bm{\tau}\tau_{xx}$ in the equation for $\dot{\tau}_{xx}$, the correct coefficient values are not obtained. Nevertheless, from Fig. 9, which shows the test simulation results, the equations obtained by the STRidge with $\alpha=1\times 10^{-3}$ and the a-Lasso with $\alpha=1\times 10^{-8}$ can well reproduce the exact solutions including the small oscillation of $\tau_{yy}$. Although the equations obtained by the STRidge and a-Lasso demonstrate similar performance in the test simulations shown in Fig. 9, the difference in predictions is quantified by their MSEs shown in Table 1. When $\alpha$ is small, for predictions of $\tau_{xx}$, $\tau_{yy}$, and $\tau_{xy}$, the STRidge outperforms the a-Lasso. The STRidge, however, generally provides sparse solutions within a narrow range of $\alpha$ values, requiring careful selection of $\alpha$.
@@
IV.3 FENE Dumbbell Model
@@
Figure 10: (a) The total number of terms and (b) the error rate of the constitutive equation obtained by Rheo-SINDy with the STRidge (black squares) and the a-Lasso (red reverse triangles) for the FENE dumbbell model. The horizontal short-dashed line in (a) indicates that the number of terms is zero. The blue circles represent the $\alpha$ values selected for test simulations.
@@
Figure 11: The constitutive equations obtained by Rheo-SINDy with the a-Lasso for the FENE dumbbell model (left) and the test simulation results under the oscillatory shear flows with $\gamma_{0}=3$ and $\omega=1$ (right upper panel) and $\gamma_{0}=4$ and $\omega=1$ (right lower panel). The bold lines show the exact solutions, and the thin solid and short-dashed lines show the results with the smaller $\alpha$ value ($\alpha=1\times 10^{-6}$) and the larger $\alpha$ value ($\alpha=3\times 10^{-4}$).
@@
Finally, we address the FENE dumbbell model. As explained in Sec. III.3, the FENE dumbbell model does not have an analytical expression of the constitutive equation. Thus, we here aim to identify an approximate constitutive equation using Rheo-SINDy.
@@
To prepare the library $\bm{\Theta}$ for the FENE dumbbell model, we utilize the physical insights obtained from the analytical expression of the FENE-P dumbbell model. We here assume the constitutive equation for the FENE-P dumbbell model is similar to that for the FENE dumbbell model. Since the FENE-P dumbbell model is a simplified version of the FENE dumbbell model, we believe that this is a reasonable assumption. Here, we note that the stress expression shown in Eq. (23) is no longer applicable to the FENE dumbbell model since the values of $h(t)$ differ for each individual dumbbell. Thus, it is invalid to obtain stress through the conformation tensor $\bm{C}$. Based on the above considerations, we decided to use the custom library presented in Eq. (33), which was also used to discover the constitutive equation for the FENE-P dumbbell model.
@@
Figure 10 compares (a) the total number of terms and (b) the error rate predicted by the STRidge and a-Lasso. Similar to the previous discussions, we obtain sparse solutions over a wide range of $\alpha$ values with the a-Lasso, whereas the STRidge gives sparse solutions only within a limited range of $\alpha$. The left table in Fig. 11 shows the equations obtained by the a-Lasso with two $\alpha$ values chosen from the viewpoints of the sparsity and error rate in the same way as Fig. 7. We note that the predictions obtained by the STRidge are inferior to those obtained by the a-Lasso shown in Fig. 11, which is discussed in Sec. S6 in the supplementary material. From the left table in Fig. 11, if $\alpha$ is appropriately chosen, the a-Lasso can give sparse equations with coefficients of reasonable (not excessively large) magnitudes. Comparing the equations for the FENE-P model obtained by the a-Lasso with $\alpha=1\times 10^{-8}$ (Fig. 8) and those for the FENE model obtained by the a-Lasso with $\alpha=1\times 10^{-6}$ (Fig. 11), the appearing terms are almost identical, which demonstrates the similarity between these models. The difference in the coefficients thus represents the difference between these models. The right panels in Fig. 11 show the test simulation results obtained by the equations shown in the left table. We found that the equations obtained with $\alpha=1\times 10^{-6}$ can reproduce well the BD simulation results outside the range of the training data within the investigated parameters, including the oscillatory behavior of $\tau_{yy}$. With the large $\alpha$ ($\alpha=3\times 10^{-4}$), the identified equation for $\tau_{yy}$ becomes ${\dot{\tau}}_{yy}=0$, which fails to reproduce the oscillatory behavior of $\tau_{yy}$. This success suggests that Rheo-SINDy with the a-Lasso is effective for discovering unknown constitutive equations. Nevertheless, we note that the equations presented in Fig. 11 may fail to predict test data significantly outside the range of the training data. Reproducing such highly nonlinear data would require the nonlinear terms dropped in Fig. 11. In this sense, the constitutive equations for the FENE dumbbell model obtained here are appropriately referred to as the approximate constitutive equations.
@@
To further investigate rheological quantities under shear flow, we conducted shear simulations under constant shear rates. Figure 12 compares the steady state shear viscosity $\eta\equiv\sigma_{xy}/\dot{\gamma}$ and first normal stress coefficient $\Psi_{1}\equiv(\sigma_{xx}-\sigma_{yy})/\dot{\gamma}^{2}$ obtained from BD simulations with those predicted by Rheo-SINDy. Although the constant shear flows considered in Fig. 12 differ from the oscillatory shear flows used to generate the training data, Rheo-SINDy can reproduce the results of BD simulations, including shear thinning, within the shear rate region of the training data and slightly beyond it, as shown in Fig. 12(a). However, in the high shear rate region, the Rheo-SINDy predictions with $\alpha=1\times 10^{-6}$ exhibit plateaus for $\eta$ and $\Psi_{1}$ that differ from the results of BD simulations. To improve the Rheo-SINDy predictions in the high shear rate region, we tested the training data with larger $\gamma_{0}$ ($\gamma_{0}=8$) while keeping the same $\omega$ values as those used to obtain Figs. 10 and 11. We note that the generated data with $\gamma_{0}=8$ show nonlinear stress responses, as shown in Sec. S7 in the supplementary material. Figure 12(b) indicates that the model found by Rheo-SINDy with the training data that includes higher shear rates successfully extrapolates continuous shear thinning to the high shear rate regime beyond the training data.
@@
Thanks to the equations obtained using Rheo-SINDy, it is possible to provide a physical interpretation with the assistance of rheological knowledge. For example, from the comparison of the equations obtained for the FENE-P dumbbell model (cf. Fig. 8) and those for the FENE dumbbell model (cf. Fig. 11), the equations for larger $\alpha$ value ($\alpha=1\times 10^{-4}$ for the FENE-P dumbbell model and $\alpha=3\times 10^{-4}$ for the FENE dumbbell model) are similar except for the coefficient values. Furthermore, the terms in these equations are the same as those for the UCM model (and thus the Hookean dumbbell model). This indicates that all of these models share the same origin based on the dumbbell model. We can also discuss the linear term of stress, which represents the relaxation of stress (see Eq. (S5) in the supplementary material). Since the relaxation time at equilibrium ($\lambda=\zeta/4h_{\rm eq}$) is taken as the unit time in this study, the coefficient of this term should be $-1$ at equilibrium (and thus for the UCM model, see Eqs. (S6)–(S8)). From Figs. 8 and 11, the coefficient of the linear term of stress with the larger $\alpha$ values is smaller than $-1$, which indicates $\lambda_{\rm sf}<\lambda_{\rm eq}$ with the subscript “sf” and “eq” standing for “shear flow” and “equiliblium”, respectively. This indicates that under shear flow, the values of spring strength for the FENE-P and FENE dumbbell models become larger than $h_{\rm eq}$, which implies the appearance of the FENE effects under flow. From this discussion, it is evident that Rheo-SINDy can provide physically interpretable constitutive equations.

### v3

@@ IV.4 FENE-P Dumbbell Model
Figure 10 shows (a) the total number of terms and (b) the error rate of the constitutive equation obtained by Rheo-SINDy with the STRidge and a-Lasso for the FENE-P dumbbell model. Similar to the previous case, the a-Lasso yields sparser solutions than the STRidge. Based on the number of terms and the error rates shown in Fig. 10, we chose two $\alpha$, one is based on the proposed criterion and the other has a smaller number of terms. Figure 11 presents the equations obtained using the chosen $\alpha$. From Fig. 11, the equations predicted by the STRidge with $\alpha=1$ and the a-Lasso with $\alpha=1\times 10^{-4}$ are almost the same; conversely, the expressions for small $\alpha$ values significantly differ between the two methods. For the STRidge with $\alpha=1\times 10^{-3}$, the identified equations are close to the correct equations (cf. Eqs. (42)–(44)). Furthermore, the coefficient values for the terms obtained correctly are close to the correct values. For the a-Lasso with $\alpha=1\times 10^{-8}$, several coefficients for the correctly obtained terms, such as $\tau_{xx}$, $\tau_{xy}\kappa_{xy}$, and $\tau_{xx}\tau_{xy}\kappa_{xy}$ in the equation for $\dot{\tau}_{xx}$, are close to the exact values, but for other several terms, such as ${\rm tr}\bm{\tau}\tau_{xx}$ in the equation for $\dot{\tau}_{xx}$, the correct coefficient values are not obtained. Nevertheless, from Fig. 12, which shows the test simulation results, the equations obtained by the STRidge with $\alpha=1\times 10^{-3}$ and the a-Lasso with $\alpha=1\times 10^{-8}$ can well reproduce the exact solutions including the small oscillation of $\tau_{yy}$. Although the equations obtained by the STRidge and a-Lasso demonstrate similar performance in the test simulations shown in Fig. 12, the difference in predictions is quantified by their MSEs shown in Table 4. When $\alpha$ is small, the STRidge outperforms the a-Lasso. The STRidge, however, generally provides sparse solutions within a narrow range of $\alpha$ values, requiring careful selection of $\alpha$. In the case of the FENE-P dumbbell model, as shown in Figs. 7 and 10, the error rate obtained by the STRidge is smaller than that obtained by the a-Lasso over the investigated $\alpha$ region. This indicates that the STRidge is superior to the a-Lasso in representing the training data. However, as shown in the next section, even with a low error rate, the performance in test simulations for the extrapolation region is not guaranteed. This indicates that the optimization method needs to be chosen based on the specific problem.
@@
IV.5 FENE Dumbbell Model
@@
Figure 13: (a) The total number of terms and (b) the error rate of the constitutive equation obtained by Rheo-SINDy with the STRidge (black squares) and the a-Lasso (red reverse triangles) for the FENE dumbbell model. The horizontal short-dashed line in (a) indicates that the number of terms is zero. The blue circles represent the $\alpha$ values selected for test simulations. Based on our criterion, the appropriate $\alpha$ values were determined as $\alpha=1\times 10^{-1}$ for the STRidge and $\alpha=1\times 10^{-6}$ for the a-Lasso.
@@
Figure 14: The constitutive equations obtained by Rheo-SINDy with the a-Lasso for the FENE dumbbell model (left) and the test simulation results under the oscillatory shear flows with $\gamma_{0}=3$ and $\omega=1$ (right upper panel) and $\gamma_{0}=4$ and $\omega=1$ (right lower panel). The black, blue, and red lines represent the $xx$-, $yy$-, and $xy$-components of the stress tensor, respectively. The bold lines show the exact solutions, and the thin solid and short-dashed lines show the results with the smaller $\alpha$ value ($\alpha=1\times 10^{-6}$), which is chosen based on our criterion, and the larger $\alpha$ value ($\alpha=3\times 10^{-4}$).
@@
We next address the FENE dumbbell model. As explained in Sec. III.3.3, the FENE dumbbell model does not have an analytical expression of the constitutive equation. Thus, we here aim to identify an approximate constitutive equation using Rheo-SINDy.
@@
To prepare the library $\bm{\Theta}$ for the FENE dumbbell model, we utilize the physical insights obtained from the analytical expression of the FENE-P dumbbell model. We assume the constitutive equation for the FENE-P dumbbell model is similar to that for the FENE dumbbell model. Since the FENE-P dumbbell model is a simplified version of the FENE dumbbell model, we believe that this is a reasonable assumption. Here, we note that the stress expression shown in Eq. (36) is no longer applicable to the FENE dumbbell model since the values of $h(t)$ differ for each individual dumbbell. Thus, it is invalid to obtain stress through the conformation tensor $\bm{C}$ using Eq. (36). Based on the above considerations, we decided to use the custom library presented in Eq. (46), which was also used to discover the constitutive equation for the FENE-P dumbbell model.
@@
Figure 13 compares (a) the total number of terms and (b) the error rate predicted by the STRidge and a-Lasso. Similar to the previous discussions, we obtain sparse solutions over a wide range of $\alpha$ values with the a-Lasso, whereas the STRidge gives sparse solutions only within a limited range of $\alpha$. The left table in Fig. 14 shows the equations obtained by the a-Lasso with two $\alpha$ values chosen in the same way as Fig. 10. The predictions obtained by the STRidge are inferior to those obtained by the a-Lasso shown in Fig. 14; therefore, we focus on the a-Lasso result below (the STRidge results are discussed in Sec. S3 in the supplementary material). From the left table in Fig. 14, if $\alpha$ is appropriately chosen, the a-Lasso can give sparse equations with coefficients of reasonable (not excessively large) magnitudes. Comparing the equations for the FENE-P model obtained by the a-Lasso with $\alpha=1\times 10^{-8}$ (Fig. 11) and those for the FENE model obtained by the a-Lasso with $\alpha=1\times 10^{-6}$ (Fig. 14), the appearing terms are almost identical, which demonstrates the similarity between these models. The difference in the coefficients thus represents the difference between these models. The right panels in Fig. 14 show the test simulation results obtained by the equations shown in the left table. With the large $\alpha$ ($\alpha=3\times 10^{-4}$), the identified equation for $\tau_{yy}$ becomes ${\dot{\tau}}_{yy}=0$, which fails to reproduce the oscillatory behavior of $\tau_{yy}$. On the contrary, the equations obtained with $\alpha=1\times 10^{-6}$ can reproduce well the BD simulation results outside the range of the training data, including the oscillatory behavior of $\tau_{yy}$.
This success suggests that Rheo-SINDy with the a-Lasso is effective for discovering unknown constitutive equations. Nevertheless, we note that the equations presented in Fig. 14 may fail to predict test data significantly outside the range of the training data. Reproducing such highly nonlinear data would require the nonlinear terms dropped in Fig. 14. In this sense, the constitutive equations for the FENE dumbbell model obtained here are appropriately referred to as the approximate constitutive equations.
@@
To further investigate rheological quantities under shear flow, we conducted shear simulations under constant shear rates. Figure 15 compares the steady state shear viscosity $\eta\equiv\sigma_{xy}/\dot{\gamma}$ and first normal stress coefficient $\Psi_{1}\equiv(\sigma_{xx}-\sigma_{yy})/\dot{\gamma}^{2}$ obtained from BD simulations with the approximate constitutive equations obtained by Rheo-SINDy. Although the constant shear flows considered in Fig. 15 differ from the oscillatory shear flows used to generate the training data, Rheo-SINDy can reproduce the results of BD simulations, including shear thinning, within the shear rate region of the training data and slightly beyond it, as shown in Fig. 15(a). However, in the high shear rate region, the Rheo-SINDy predictions with $\alpha=1\times 10^{-6}$ exhibit plateaus for $\eta$ and $\Psi_{1}$ that differ from the results of BD simulations. To improve the Rheo-SINDy predictions in the high shear rate region, we tested the training data with larger $\gamma_{0}$ ($\gamma_{0}=8$) while keeping the same $\omega$ values as those used to obtain Figs. 13 and 14. We note that the generated data with $\gamma_{0}=8$ show nonlinear stress responses, as shown in Sec. S4 in the supplementary material. Figure 15(b) indicates that the model found by Rheo-SINDy with the training data that includes higher shear rates successfully extrapolates continuous shear thinning to the high shear rate regime beyond the training data.
Figure 16: (a) Transient state shear viscosity $\eta^{+}$ and (b) first normal stress coefficient $\Psi_{1}^{+}$ under the steady shear flow with $\dot{\gamma}=1$ (solid lines) and $10$ (dotted lines) for the FENE-P (black) and FENE (red) dumbbell models. The thin and bold lines were obtained by Rheo-SINDy and by the equations for the FENE-P and FENE dumbbell models, respectively. The Rheo-SINDy regressions were conducted for the training data using the same parameters ($\gamma_{0}=8$ and $\omega\in\{0.1,0.2,\ldots,1\}$) for both the FENE-P and FENE dumbbell models. While the Rheo-SINDy results for the FENE-P dumbbell model were obtained by the STRidge with $\alpha=1\times 10^{-4}$, those for the FENE dumbbell model were obtained by the a-Lasso with $\alpha=1\times 10^{-5}$, which is the same parameter as that in Fig. 15(b).
It would be interesting to compare the results obtained by Rheo-SINDy for the FENE-P and FENE dumbbell models. Figure 16 indicates the transient shear viscosity $\eta^{+}$ and first normal stress coefficient $\Psi_{1}^{+}$ obtained by Rheo-SINDy and those exact solutions for the FENE-P and FENE dumbbell models. We note that, for comparison, the Rheo-SINDy regressions were conducted using the training data with the same parameters ($\gamma_{0}=8$ and $\omega\in\{0.1,0.2,\ldots,1\}$) for both models. For the FENE-P dumbbell model, the regression for the stress expression (cf. Eqs. (42)–(44)) was performed. From the success of the STRidge in this case (cf. Fig. 10), we used the STRidge with $\alpha=1\times 10^{-4}$. The Rheo-SINDy results for the FENE dumbbell model were obtained using the a-Lasso with $\alpha=1\times 10^{-5}$, which is the same parameter as that determined in Fig. 15(b). Figure 16 shows that Rheo-SINDy for the FENE-P dumbbell model accurately reproduces the transient behavior shown in the exact solutions. However, while Rheo-SINDy for the FENE dumbbell model successfully reproduces the steady state values of $\eta^{+}$ and $\Psi_{1}^{+}$, as expected from Fig. 15(b), it fails to describe the stress overshoot, especially for the large $\dot{\gamma}$. Although the objective of this study is not to develop a sophisticated constitutive model that approximates the FENE dumbbell model, this discrepancy can be addressed through three possible strategies: (1) adding training data obtained by steady shear flow with high shear rate, (2) using the conformation tensor $\bm{C}$ as a latent variable, and (3) incorporating higher-order terms into the library for regression. For point (1), it has been reported that including training data obtained from constant shear flow with high shear rate is effective Miyamoto, Molina, and Taniguchi (2023). This enables the development of regression models that can predict transient behavior, even for more complex models. For point (2), the conformation tensor often allows for simpler descriptions, as shown for the FENE-P dumbbell model. It might be effective to use $\bm{C}$ to obtain regression equations in the FENE dumbbell model. This case will require an additional regression to relate $\bm{C}$ and $\bm{\tau}$, replacing Eq. (36) for the FENE-P dumbbell model. For point (3), Du and coworkers have reported that the approximated model with high-order terms achieves high predictive performance Du, Liu, and Yu (2005). Following this study, including higher-order terms in the library can enhance the representational ability of Rheo-SINDy. These considerations are interesting subjects for future studies.
@@
Thanks to the equations obtained using Rheo-SINDy, it is possible to provide a physical interpretation with the assistance of rheological knowledge. For example, from the comparison of the equations obtained for the FENE-P dumbbell model (cf. Fig. 11) and those for the FENE dumbbell model (cf. Fig. 14), the equations for larger $\alpha$ value ($\alpha=1\times 10^{-4}$ for the FENE-P dumbbell model and $\alpha=3\times 10^{-4}$ for the FENE dumbbell model) are similar except for the coefficient values. We note that these $\alpha$ values are not selected based on our criteria. However, to simply analyze the representation obtained by Rheo-SINDy, we consider the sparser solution here. The terms in these equations are the same as those for the UCM model (and thus the Hookean dumbbell model). This indicates that all of these models share the same origin based on the dumbbell model. We can also discuss the linear term of stress, which represents the relaxation of stress (see Eq. (18)). Since the relaxation time at equilibrium ($\lambda=\zeta/4h_{\rm eq}$) is taken as the unit time in this study, the coefficient of this term should be $-1$ at equilibrium (and thus for the UCM model, see Eqs. (19)–(21)). From Figs. 11 and 14, the coefficient of the linear term of stress with the larger $\alpha$ values is smaller than $-1$, which indicates $\lambda_{\rm sf}<\lambda_{\rm eq}$ with the subscript “sf” and “eq” standing for “shear flow” and “equiliblium”, respectively. This indicates that under shear flow, the values of spring strength for the FENE-P and FENE dumbbell models become larger than $h_{\rm eq}$, which implies the appearance of the FENE effects under flow. From this discussion, it is evident that Rheo-SINDy can provide physically interpretable constitutive equations.
Figure 17: The total number of terms and the error rate of the constitutive equations obtained by Rheo-SINDy with the STRidge (black squares) and the a-Lasso (red reverse triangles) for (a) ${\bm{\tau}}^{\ast}$ and for (b) $\theta^{\ast}_{\bm{\tau}}$. The blue circles represent the appropriate $\alpha$ values. The horizontal lines in the top figures indicate the correct number of terms. (c) The constitutive equations obtained by Rheo-SINDy with the STRidge and the a-Lasso.
IV.6 Material Frame Indifference
Finally, we show a method to obtain constitutive models that address the principle of material objectivity. This principle requires that constitutive equations are independent of rotation. However, the results obtained by the Rheo-SINDy regression method presented thus far do not explicitly satisfy the rotational indifference.
To account for this indifference, Young and coworkers demonstrated that it is effective to separate the rotational components within the velocity gradient tensor from constitutive equations Young et al. (2023). In this approach, the velocity gradient tensor $\bm{\kappa}$ in the laboratory frame (LF) is decomposed into the deformation rate tensor $\bm{D}$ and the rotation rate tensor $\bm{W}$ as $\bm{\kappa}=\bm{D}+\bm{W}$, where $\bm{D}=(\bm{\kappa}+\bm{\kappa}^{\rm T})/2$ and $\bm{W}=(\bm{\kappa}-\bm{\kappa}^{\rm T})/2$. We then consider a constitutive model in the reference frame (RF) rotated by $\bm{W}$. Here, we note that the constitutive model in the RF should depend only on $\bm{D}$. We can compute the rotation matrix $\bm{R}(\theta_{\bm{W}})$ around an axis that relates the LF and RF as
$\frac{{\rm d}{\bm{R}}(\theta_{\bm{W}})}{{\rm d}t}={\bm{W}}\cdot{\bm{R}}(\theta_{\bm{W}}),$ (47)
where $\theta_{\bm{W}}$ is the rotation angle. Using the rotation matrix $\bm{R}(\theta_{\bm{W}})$, the stress tensor $\bm{\tau}$ and the strain rate tensor $\bm{D}$ are respectively transformed into $\bm{\tau}^{\prime}=\bm{R}^{\rm T}(\theta_{\bm{W}})\bm{\tau}\bm{R}(\theta_{\bm{W}})$ and $\bm{D}^{\prime}=\bm{R}^{\rm T}(\theta_{\bm{W}})\bm{D}\bm{R}(\theta_{\bm{W}})$, where prime denotes quantities in the RF. We note that Young and coworkers demonstrated that the constitutive equation of the UCM model in the RF can be written similarly to that in the LF using $\bm{\tau}^{\prime}$ and $\bm{D}^{\prime}$ Young et al. (2023).
Based on the study of Young and coworkers Young et al. (2023), we here propose a more rigorous approach to account for material objectivity. Note that Lennon and coworkers proposed another method to employ prescribed tensor terms satisfying the frame indifference Lennon, McKinley, and Swan (2023). Since the current Rheo-SINDy method does not address constraints of inter-coefficients for tensorial terms, the identified equations should be constructed using invariants of the stress and strain rate tensors. In this approach, $\bm{\tau}^{\prime}$ and $\bm{D}^{\prime}$ in the RF are diagonalized using rotation matrices around an axis $\bm{P}(\theta^{\ast}_{\bm{\tau}})$ and $\bm{P}(\theta^{\ast}_{\bm{D}})$ as $\bm{\tau}^{\ast}=\bm{P}^{\rm T}(\theta^{\ast}_{\bm{\tau}})\bm{\tau}^{\prime}\bm{P}(\theta^{\ast}_{\bm{\tau}})$ and $\bm{D}^{\ast}=\bm{P}^{\rm T}(\theta^{\ast}_{\bm{D}})\bm{\tau}^{\prime}\bm{P}(\theta^{\ast}_{\bm{D}})$, respectively. Here, $\theta^{\ast}_{\bm{\tau}}$ and $\theta^{\ast}_{\bm{D}}$ are the rotation angles for $\bm{\tau}^{\ast}$ and $\bm{D}^{\ast}$, respectively. With $\bm{\tau}^{\ast}_{\bm{\tau}}$, $\bm{\tau}^{\ast}_{\bm{D}}$, and $\phi^{\ast}=\theta_{\bm{\tau}}^{\ast}-\theta_{\bm{D}}^{\ast}$, the constitutive equations are expressed as follows:
$\dot{\bm{\tau}}^{\ast}=\bm{f}_{{\bm{\tau}}^{\ast}}({\bm{\tau}}^{\ast},{\bm{D}}^{\ast},\phi^{\ast})$ (48)
$\dot{\theta}_{\bm{\tau}}^{\ast}=f_{{\theta}_{\bm{\tau}}^{\ast}}({\bm{\tau}}^{\ast},{\bm{D}}^{\ast},\phi^{\ast})$ (49)
Here, $\phi^{\ast}$ indicates the adjustment for using rotation matrices with different rotation angles to obtain $\bm{\tau}^{\ast}$ and $\bm{D}^{\ast}$. This type of constitutive equations can cover constitutive relations under deformations in a plane. In the supplementary material (Sec. S5), we demonstrate the derivation of constitutive equations for $\bm{\tau}^{\ast}$ and $\bm{\theta}^{\ast}_{\bm{\tau}}$ in the UCM and Giesekus models. Hereafter, we show that Rheo-SINDy can give Eqs. (48) and (49) for the Giesekus model. From $\bm{\tau}^{\ast}$ and $\theta^{\ast}_{\bm{\tau}}$, $\bm{\tau}$ in LF can be determined by the relationship $\bm{\tau}=\bm{P}(\theta_{\bm{W}})\bm{P}(\theta^{\ast}_{\bm{\tau}})\bm{\tau}^{\ast}\bm{P}^{\rm T}(\theta^{\ast}_{\bm{\tau}})\bm{P}^{\rm T}(\theta_{\bm{W}})$.
Using the same training data shown in Sec. III.2, we here conduct the Rheo-SINDy regressions. In the Giesekus model, the constitutive equations for $\bm{\tau}^{\ast}$ and $\theta^{\ast}_{\bm{\tau}}$ are as follows (see the supplementary material for the detailed derivations):
$\dot{\tau}^{\ast}_{xx}$ $=+2\tau^{\ast}_{xx}D_{xx}^{\ast}\cos 2\phi^{\ast}-{\tau}^{\ast}_{xx}-\alpha_{\rm G}({\tau}^{\ast}_{xx})^{2}+2D_{xx}^{\ast}\cos 2\phi^{\ast}$ (50)
$\dot{\tau}^{\ast}_{yy}$ $=-2\tau^{\ast}_{yy}D_{xx}^{\ast}\cos 2\phi^{\ast}-{\tau}^{\ast}_{yy}-\alpha_{\rm G}({\tau}^{\ast}_{yy})^{2}-2D_{xx}^{\ast}\cos 2\phi^{\ast}$ (51)
$\dot{\tau}^{\ast}_{zz}$ $=-{\tau}^{\ast}_{zz}-\alpha_{\rm G}({\tau}^{\ast}_{zz})^{2}=0$ (52)
$\dot{\theta}_{\bm{\tau}}^{\ast}$ $=-\left(\frac{\tau^{\ast}_{xx}}{\tau^{\ast}_{xx}-\tau^{\ast}_{yy}}+\frac{\tau^{\ast}_{yy}}{\tau^{\ast}_{xx}-\tau^{\ast}_{yy}}+\frac{2}{\tau^{\ast}_{xx}-\tau^{\ast}_{yy}}\right)D_{xx}^{\ast}\sin 2\phi^{\ast}$ (53)
Similarly to Eqs. (23)–(26), Eqs. (50)–(53) are nondimensionalized using $\lambda$ and $G$. We prepared polynomial libraries up to second order consisting of $\{\tau^{\ast}_{xx},\tau^{\ast}_{yy},\tau^{\ast}_{zz},D^{\ast}_{xx}\sin(2\phi^{\ast}),D^{\ast}_{xx}\cos(2\phi^{\ast})\}$ for the regression of $\bm{\tau}^{\ast}$, and $\{1/\Delta,\tau^{\ast}_{xx}/\Delta,\tau^{\ast}_{yy}/\Delta,\tau^{\ast}_{zz}/\Delta,D^{\ast}_{xx}\sin(2\phi^{\ast}),D^{\ast}_{xx}\cos(2\phi^{\ast})\}$ with the shorthand $\Delta=\tau^{\ast}_{xx}-\tau^{\ast}_{yy}$ for the regression of $\theta^{\ast}_{\bm{\tau}}$. Thus, there were $N_{\bm{\Theta}}=21$ and $N_{\bm{\Theta}}=28$ candidate terms for each component of $\bm{\tau}^{\ast}$ and $\bm{\theta}^{\ast}_{\bm{\tau}}$, respectively. Based on dimensional consideration, it is natural to create the library for the regression of $\theta^{\ast}_{\bm{\tau}}$ using terms of $\bm{\tau}^{\ast}/\Delta$.
Figure 17 indicates the Rheo-SINDy results for $\bm{\tau}^{\ast}$ and $\theta^{\ast}_{\bm{\tau}}$. The trends in the number of terms and error rate shown in Figs. 17(a) and (b) are the same as those in the regressions with Rheo-SINDy at the LF (cf. Fig. 4). The lower panel in Fig. 17(a) shows that the a-Lasso has a larger error rate than the STRidge, reflecting the inaccuracies in the regression, especially in $\tau^{\ast}_{xx}$, as shown in Fig. 17(c). Nevertheless, in the regressions of $\theta^{\ast}_{\bm{\tau}}$, the STRidge and a-Lasso have similar error rates, which indicates that these methods have comparable performance in this case. We note that the expression of $\dot{\theta}^{\ast}_{\bm{\tau}}$ obtained by the a-Lasso, $\dot{\theta}^{\ast}_{\bm{\tau}}\simeq-(1+2\tau^{\ast}_{yy}/\Delta+2/\Delta)D^{\ast}_{xx}\sin 2\phi^{\ast}$, reduces the correct equation shown in Eq. (53). The Rheo-SINDy regressions for $\bm{\tau}^{\ast}$ using STRidge have small error rates, but the obtained expressions shown in Fig. 17(c) differ from Eqs. (50) and (51). This is because the Giesekus model with $\alpha_{\rm G}=0.5$ is equivalent to the Leonov model Larson (1988). Interestingly, by conducting the diagonalization, the equivalent mathematical expression becomes evident through the Rheo-SINDy regression.
Through the discussion in this section, we show that a constitutive equation that takes into account material objectivity can be obtained from Rheo-SINDy. This indicates that the technical preparations are now in place to precisely handle more complex models using Rheo-SINDy.